A copper concentrate (CuFeS2 + FeS2 + gangue) of composition (wt%) Cu = 6%, S = 35% is
roasted with (20+RN*)% excess air. Copper in the roasted product is as Cu2S and all the iron is
converted to Fe2O3. Assuming that the gangue does not contain any Cu, Fe and S, Calculate for 1
ton of concentrate i) amount of minerals and gangue ii) volume of air supplied for roasting at STP.
iii) Volume and composition of the evolved gas (flue) after roasting [atomic weight of Cu=64, Fe=
56, S=32, O=16, N=14]
1
Expert's answer
2020-11-23T06:49:52-0500

(i)Let "\omega"(CuFeS2)=c and "\omega"(FeS2)=f

c*64/(64+56+32*2) = 0.06

c*32*2/(64+56+32*2)+f*32*2/(56+32*2) = 0.35


c*0.348 = 0.06

c*0.348+f*0.533 = 0.35


c=0.172; f=0.334

"\omega"(CuFeS2)=0.172; "\omega"(FeS2)=0.344; "\omega"(gangue)=0.484

(ii) excess of air=20%; "\chi" [mole fraction](O2 in air)=0.21; "\chi" (N2 in air)=0.79

4CuFeS2+9O2=2Cu2S+2Fe2O3+6SO2

4FeS2+11O2=2Fe2O3+8SO2

n(CuFeS2)=1000000*0.172/(64+56+32*2)=935 mol

n(FeS2)=1000000*0.344/(56+32*2)=2867 mol

n(O2, theor)=935/4*9+2867/4*11=9988 mol

V(air)=9988/0.21*1.20*22.4/1000=1278 m3

(iii)

V(O2)=9988*0.2*22.4/1000=44.7 m3

V(N2)=1278*0.79=1010 m3

V(SO2)=(935/4*6+2867*2)*22.4/1000=160 m3

V(flue)=44.7+1010+160=1215 m3

"\chi" (N2)=0.831; "\chi" (SO2)=0.132; "\chi" (O2)=0.037.